OpenStudy (anonymous):

2x-y=3,9x-4y=15 how would i do this in the substition method

OpenStudy (neo92):

make y or x the subject in one eqn and substitute it in the other. \[2x-y=3\] when y is made the subject \[y=2x-3\]Now substitute (2x-3) for y in the second eqn

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

y=9x-4

OpenStudy (anonymous):

is that right

OpenStudy (neo92):

nope.. from 1st eq we got y = 2x - 3 correct?

OpenStudy (anonymous):

yes

OpenStudy (neo92):

now if we make y the subject in the 2nd eq, we get 4y = 9x - 15 got it?

OpenStudy (anonymous):

yes

OpenStudy (neo92):

now substitute the value we got for y frm 1 in 2. that is 4(2x-3) = 9x-15

OpenStudy (neo92):

got it? should i proceed or can u do it by urself?

OpenStudy (anonymous):

no keep going or was that it

OpenStudy (anonymous):

is there more to it

OpenStudy (anonymous):

hello

OpenStudy (neo92):

sorry 4 da l8 amigo.. rest is quite simple.. just solve the final eqn we got to find x 4(2x-3) = 9x-15 8x-12 = 9x-15 get it now?

OpenStudy (anonymous):

is the answer

OpenStudy (neo92):

x = 3 now u can subtitute this value in any of the initial eqns and find y.. get it?

OpenStudy (anonymous):

is there more to it

OpenStudy (neo92):

u hav to find both x n y.. now we got x = 3 can u find y?

OpenStudy (anonymous):

not sure how to do that

OpenStudy (neo92):

from 1st eq we got y = 2x -3 Now we know x = 3 so what is y?

OpenStudy (anonymous):

3

OpenStudy (neo92):

yep.. :D

OpenStudy (anonymous):

so it s x=3 and y=3

OpenStudy (neo92):

yep bro..

OpenStudy (anonymous):

thanks